Solve the equation: log(base2)2-log(base)8=2
Solve the simultaneous equations:a^3+b^3 and a+b=2

Answer by stanbon(75887)

Solve the equation:
log(base2)2-log(base)8=2
-----
1 - (1/3)
-----
= 2/3
=========================
Solve the simultaneous equations:a^3+b^3 and a+b=2
Factor:
a^3+b^3 = (a+b)(a^2 - ab + b^2)
-----
= 2(a^2-ab+b^2)
==================
Cheers,
Stan H.
==================